There were 2 mens in a garden.They make a circle of diameter 28cm and with radius 14cm.Find the area of the circle?

Grade:8

3 Answers

Saurabh Koranglekar
4 years ago
Dear student

Area of circle is pi * r*r

Pi = 22/7
R = 7

Area is 184 cm sq

Regards


Vikas TU
4 years ago
Dear student 
Above ans is incorrect 
Diameter = 28 cm
Radius = diameter/2
          = 28/2
          = 14 cm
Area of circle = Π r ²
                   =  ( 22/7 )  x (14)²
                   = ( 22/7 )  x 14 x 14
                   = 22 x 2 x 14
                   = 44 x 14
                   = 616 cm²
 
Yashraj
4 years ago
radius =14cm
pi =22/7
area = pi * R2
area =22/7*169=
=616 cm2
hope good luck.................................................